OpenStudy (anonymous):

I really need to be sure I get a 100 on this:( What is the axis of symmetry for the graph of y=2x^2-8x+2? What is the vertex for the graph of Y-2=-(x+3)^2? What are the x-intercepts for the graph of y=3(x-2)(x+6)? PLEASE HELP!

OpenStudy (triciaal):

y=2x^2-8x+2 do you know how to complete the square? put in vertex form vertex form y = a(x-h)^2 + k vertex = (h, k) axis of symmetry x = h h= -b/(2a)

OpenStudy (triciaal):

What is the vertex for the graph of Y-2=-(x+3)^2? rewrite as y = vertex (h,k)

OpenStudy (triciaal):

What are the x-intercepts for the graph of y=3(x-2)(x+6)? means what are the x values when y = 0 when y = 0 means 3 =0 or (x-2) = 0 or (x + 6) = 0

OpenStudy (triciaal):

any questions about my response?

OpenStudy (anonymous):

Thank you! I'm just trying to figure it all out